OpenStudy (anonymous):

The side of a large tent is in the shape of an isosceles triangle whose area is 54 ft² and whose base is 6 feet shorter than twice its height. Find the height and the base of the tent (height by base).

OpenStudy (tad1):

b = 2h - 6 A = (1/2)bh= (1/2) (2h_6)h=54 Solve for h back substitute for b

OpenStudy (anonymous):

Thank You So Much I Got 9 ft by 12 ft